# Chrono[Uncaptioned image]: A Simple Blueprint for Representing Time in MLLMs

### v5

@@ 3.2 Training and Zero-Shot Chrono Setup
In this section, we describe two different options to apply the Chrono blueprint. We can either train our generative MLLM or perform zero-shot adaptation on the task of moment retrieval. We cast the traditional computer vision problem of moment retrieval as an open-ended natural language sequence-to-sequence task.
@@
Training Setup. To train our model, we optimize for the standard maximum likelihood objective. To further optimize training and best leverage the limited video data available, we implement parameter-efficient finetuning via LoRA (Hu et al., 2022) and randomly sample video frames. This allows us to only train a fraction of the parameters of the MLLM. Additionally, by randomly sampling the frames, we optimize those weights by analyzing different frames and timestamps of the same videos each epoch. For further details, see Section B.2.2.
@@
For our experiments specifically, we leverage an encoder-decoder style MLLM, BLIP-2 (Li et al., 2023). We denote BLIP-2’s frozen image encoder combined with Q-Former as our general-purpose frame encoder. The frame encoder is separately applied to each of the $F$ sub-sampled frames, generating the frame embeddings and projecting them into the language space, yielding a shared embedding space. Prior works (Wang et al., 2024; Ren et al., 2023; Yang et al., 2023; Zhang et al., 2023b) incorporate further trainable transformer-based modules to generate temporally and contextually correlated frame- or video-level embeddings. In contrast, we solely leverage the self- and cross-attention mechanisms of our LLM to learn the temporal and contextual relationships between frames. We coin this instantiation of the Chrono blueprint Chrono-BLIP.
@@
Zero-Shot Setup. For zero-shot adaptation, we can apply our Chrono framework to an MLLM with instruction-following abilities by providing it with the video frames interleaved with their respective timestamps, the video duration, user query, and instruction prompt, just as in the training setup. The difference is that we modify the instruction prompt so that our MLLM backbone understands and solves the desired task and returns the desired format.
For our experiments, we want to demonstrate the ability of the Chrono blueprint on another MLLM architecture, the decoder-only architecture, and therefore select GPT-4o (OpenAI, 2024a) as our MLLM backbone of choice. We call this instantiation of our framework Chrono-GPT and provide the full prompt used in Appendix Figure 4.
@@
Table 1: Ablation studies. We study (a) prompt components and (b) timestamp design for both our finetuned Chrono-BLIP and zero-shot Chrono-GPT. For Chrono-BLIP, we ablate on the Charades-STA validation set, and for Chrono-GPT, on the QVH validation set. D: Video Duration, T: Frame Timestamps, Rep: Representation, Rel: Relative, Abs: Absolute, Prec: Precision, Dec: Decimal, Int: Integer, Inter: Interleaved.
@@
4 Experiments
This section describes the effectiveness of our design choices and compares our method to the state-of-the-art. We validate Chrono on the three most widely used video moment retrieval (MR) datasets Charades-STA (Gao et al., 2017), QVHighlights (Lei et al., 2021a), and ActivityNet Captions (Krishna et al., 2017), and extend our framework to the task of grounded video questions answering on the NExT-GQA (Xiao et al., 2023) benchmark. We present and analyze the experiments in this section, with further details on our experimental setup in Appendix B.
4.1 Ablation Studies
By default, our Chrono models for Charades-STA (QVHighlights/ ActivityNet Captions) leverage 20 (60) visual frames, their associated timestamps, video duration, a query, and a task prompt. If not stated otherwise, we represent time as simple text tokens of whole seconds and interleave these with frame tokens. In the following, we ablate the impact of these design choices on the downstream moment retrieval performance by reporting results for Chrono-BLIP on our Charades-STA validation set and for Chrono-GPT on the QVHighlights validation set.
4.1.1 Design of sequence
In Table 1 (a), we analyze the effectiveness of each part of the multimodal input sequence of our MLLM by cumulatively adding each component, starting with only providing the frames. The query is always part of the sequence. For Chrono-BLIP, adding the video duration, in addition to the visual inputs, improves the downstream performance significantly (row 2 vs. row 1). This shows the importance of providing the model a point of reference to infer the temporal location of each frame and demonstrates the ability of the MLLM to learn this association. For Chrono-GPT, on the other hand, does not benefit as strongly from the additional information as our fine-tuned BLIP-2 model. Again, note that with a fixed number of uniformly sampled frames and the video duration, the model has all the information necessary to compute at which timestamp each frame was sampled. Our experiment shows that GPT-4o appears not to be able to leverage this information.
Yet, providing the specific timestamps at which each frame was sampled in an interleaved manner (as illustrated in Figure 2) improves the model performance of both our models by a significant margin (rows (3, 4) vs. rows (1,2)). This demonstrates the relevance of designing the input context to provide as much relevant information as possible while still not relying on additional computation, such as generating and leveraging transcripts. Providing all timestamps along with the video duration yields the best performance (row 4 vs. 3) for both models.

### v6

@@ 3.2 Training and Zero-Shot Chrono Setup
In this section, we describe two different options to apply the Chrono blueprint. We can either train the generative MLLM or perform zero-shot adaptation on the task of moment retrieval. We cast the traditional computer vision problem of moment retrieval as an open-ended natural language sequence-to-sequence task.
3.2.1 Training Setup
@@
To train the model, we optimize for the standard maximum likelihood objective. In order to make finetuning less compute intensive, we employ parameter-efficient finetuning via LoRA [hu2021lora] and randomly sample video frames. This allows us to only train a fraction of the parameters of the MLLM. We use a rank of 8 and apply it to all linear layers in the large language model, yielding about 19 million trainable parameters for Chrono-BLIP. Additionally, by randomly sampling the frames, we optimize those weights by analyzing different frames and timestamps of the same videos each epoch. In Section 4.6.4 and Section 4.6.5, we study the effect of the number of trainable parameters and the number of frames, respectively. We provide more details in Appendix A.
To demonstrate the applicability of the Chrono[Uncaptioned image] blueprint to different model architectures, we explore both finetuning BLIP-2 [li2023blip2] and Qwen2.5-VL [Qwen2.5-VL]. BLIP-2 [li2023blip2] is an encoder-decoder style MLLM, whereas Qwen2.5-VL [Qwen2.5-VL] is a decoder-only model.
@@
Finetuning BLIP-2. For BLIP-2, we use its frozen image encoder combined with Q-Former as the general-purpose frame encoder. The frame encoder is separately applied to each of the $F$ sub-sampled frames, generating the frame embeddings and projecting them into the language space, yielding a shared embedding space. Prior works [Wang2024InternVideo2SF, Ren2023TimeChatAT, yang2023vid2seq, zhang2023videollama] incorporate further trainable transformer-based modules to generate temporally and contextually correlated frame- or video-level embeddings. In contrast, we solely leverage the self- and cross-attention mechanisms of the LLM to learn the temporal and contextual relationships between frames. We coin this instantiation of the Chrono blueprint Chrono-BLIP.
Finetuning Qwen2.5-VL. We explore finetuning two models of the Qwen2.5-VL [Qwen2.5-VL] family, the 3 and 7 billion parameter versions. Qwen2.5-VL jointly finetunes a ViT encoder and an LLM decoder. The ViT encoder applies full attention across all patches of all images in some layers if a succession of images is embedded. However, if two images are separated by some text, like in interleaved timestamp setting, the images are encoded by the ViT independently. We denote applying the Chrono blueprint to Qwen2.5-VL Chrono-Qwen. Notably, we also show that Qwen2.5-VL can be used in a zero-shot manner for moment retrieval, although its performance is significantly lower in comparison to when finetuned.
3.2.2 Zero-Shot Setup
@@
For zero-shot adaptation, we can apply the Chrono framework to an MLLM with instruction-following abilities by providing it with the video frames interleaved with their respective timestamps, the video duration, user query, and instruction prompt, just as in the training setup. We observe that additional format prompting helps the MLLM in this zero-shot setting.
For the zero-shot experiments, we aim to demonstrate the ability of the Chrono blueprint on a model which we do not finetune for moment retrieval. Most zero-shot experiments use GPT-4o [GPT4o] since it is a stronger model.
We call this instantiation of the framework Chrono-GPT and provide the full prompt used in Appendix Figure 4. We also demonstrate the advantage of using the Chrono blueprint in a zero-shot setup using Qwen2.5-VL [Qwen2.5-VL] models in Section 4.6.3.
4 Experiments
This section describes the effectiveness of the design choices in this work and compares the proposed method to the state-of-the-art. We validate Chrono on the three most widely used video moment retrieval (MR) datasets Charades-STA [gao2017charadesSTA], QVHighlights [lei2021qvhighlights], and ActivityNet Captions [krishna2017ActivityNetCaptions], and extend the framework to the task of grounded video question answering on the NExT-GQA [Xiao2023NExT-QGA] benchmark.
We present and analyze the experiments in this section, as well as introduce the benchmarks used and metrics studied. We first compare to SOTA methods, then present qualitative results, and finally ablate key design choices.
4.1 Benchmarks
4.1.1 Moment Retrieval
Charades-STA [gao2017charadesSTA] includes 9,848 videos with an average duration of 30.6 seconds. The dataset contains 16,128 annotations with an average moment length of 8.1 seconds and an average query length of 7.22 words. The dataset is originally divided into two splits: training (12,408) and test (3,720). To avoid overfitting on the test set during training, we designate a part of the training set as a new validation set. The videos used in the validation set are not contained in the training set. The new dataset split consists of a split for training, validation, and testing, with 11,166, 1,242, and 3,720 annotations, respectively. For a fair comparison, after completing the ablations, we train the final model on the original training set and report numbers on the test set. We will share this split for reproducibility.
QVHighlights [lei2021qvhighlights] is one of the most recent benchmarks and contains 10,148 videos with a duration of 150 seconds. The videos were cropped out of YouTube videos, and each video is annotated with at least one query with an average length of 11.3 words describing the relevant moment. The target windows have an average length of 24.6 seconds. The dataset is split into training, validation, and test sets with 7,218, 1,150, and 1,542 queries, respectively. This benchmark is challenging because one query can be associated with multiple moments in a video. The test set targets are withheld and thus guarantee a fair benchmark. The evaluation for the test split can only be measured through submitting the prediction to the evaluation server [^2] Eval server: https://codalab.lisn.upsaclay.fr/competitions/6937.
ActivityNet Captions [krishna2017ActivityNetCaptions] contains 20,000 videos with an average duration of 2 minutes. The dataset contains 72,000 segments, each human-annotated with a caption that includes, on average, 13.5 words. The dataset is divided into three splits, train (37,421), val_1 (17,505), and val_2 (17,031). Following [yan2023unloc], we use the train split for training, val_1 for validation, and val_2 for testing.
4.1.2 Temporally Grounded Question Answering
NExT-GQA [Xiao2023NExT-QGA] extends the NExT-QA [Xiao2021NExTQA] benchmark by providing temporal grounding for the moments in the video, that are relevant for answering the question for the validation and test sets, making this a weakly-supervised dataset. The dataset contains the original training split of NExT-QA, which includes 34,132 samples. The validation set contains 3,358 questions with 3,931, meaning a sample can potentially have multiple relevant moments, but where predicting any single one of those moments is a correct prediction, i.e., one does not have to predict all relevant moments. The test set has 5,553 questions with 6,600 segments. The segments have an average duration of 7.3 and 6.7 seconds for the validation and test set, respectively.
4.2 Metrics
The most commonly used metrics for MR are Recall@K and mean average precision (mAP) computed under different Intersection over Union (IoU) thresholds. The Recall@K metric is defined as the percentage of the top-K predicted segments having a larger temporal IoU than the threshold with a ground truth segment. Following the recent development [lee2023bamdetr, lei2021qvhighlights, moon2023CQ_DETR, moon2023QD-DETR], we report the more challenging $R1@0.5$ and $R1@0.7$ scores, which correspond to the $Recall@1$ scores at IoU thresholds of $0.5$ and $0.7$, respectively, and the mean IoU (mIoU) score. As in these prior works, for MR on QVHighlights, we report the mAP score at IoU thresholds of $0.5$ and $0.75$ and the average mAP, respectively.
NExT-GQA [Xiao2023NExT-QGA] introduces the metrics (i) Intersection over Prediction (IoP), which is defined as the ratio between the length of the temporal intersection between the predicted relevant moment and ground truth and the length of the predicted moment itself, and (ii) Accuracy@GQA (A@GQA) which is the accuracy of correctly answered questions that have a grounding score of $IoP>0.5$. The IoP loosens the common IoU metric by only considering the intersection over the length of the prediction instead of over the union. This metric encourages very short-moment predictions and is not concerned with a precise coverage of the entire relevant moment. Therefore, following prior work, we report mean IoU (mIoU) together with IoP in NExT-GQA.
4.3 Comparison to the SOTA in Moment Retrieval
TABLE I: Comparison to state-of-the-art methods in moment retrieval (left) and grounded video question answering (right). Bold numbers indicate best result for a given benchmark and metric among all methods.
(a) Moment Retrieval on QVHighlights, Charades-STA, and ActivityNet Captions test sets. For QVHighlights, indicates validation set was used.
QVHighlights Charades-STA ActivityNet
Method R1@.5 R1@.7 mAP mAP@.5 mAP@.75 R1@.5 R1@.7 R1@.5 R1@.7
Finetuned for Moment Retrieval
VLG 46.30 29.80
SeViLa 54.50 36.50
CG-DETR 65.43 48.38 42.90 64.51 42.77 58.44 36.34
UnLoc-L 60.80 38.40 48.30 30.20
EaTR 61.36 45.79 41.74 61.86 41.91 68.47 44.92
InternVideo2-6B 71.42 56.45 49.24 70.03 48.95
BLIP-2 69.10 46.52 37.87 60.68 38.93 43.33 32.60 25.84 9.72
Chrono-Qwen[Uncaptioned image] 73.48 55.10 42.83 63.74 44.47 67.98 45.22 51.25 32.79
Zero-Shot
GPT-4o 5.49 2.17 1.10 2.81 0.76 7.02 2.12 7.20 2.80
(b) Grounded Video Question Answering on NExT-GQA test set. First four models are known to be explicitly finetuned on NExT-QA. See discussion in Section $4.$
Method mIoU IoP@.5 A@GQA A@QA
Finetuned on Dataset
SeViLA [yu2023SeViLA] 21.7 22.9 16.6 68.1
FrozenBiLM [Xiao2023NExT-QGA] 9.6 23.7 17.5 70.8
Chrono-BLIP[Uncaptioned image] 28.7 24.6 19.4 73.9
Zero-Shot
Uniform 60F 21.1 21.1 8.7 7.1
LLoVi [Zhang2023LLoVi] 21.5 38.0 26.8 -
DeVi [Qin2024DeVi] 22.3 37.9 28.0 71.6
Chrono-GPT[Uncaptioned image] 36.5 50.8 42.1 79.3
We compare Chrono to the state-of-the-art (SOTA) approaches. Specifically, we compare Chrono-BLIP and Chrono-Qwen to other finetuned baselines and Chrono-GPT to zero-shot baselines.
Firstly, in Table I(a), we include a comparison to a vanilla BLIP-2 baseline (frames-only, without the Chrono sequence design), where we finetune and evaluate it on each dataset, respectively. As expected, the Chrono-BLIP model significantly outperforms this baseline. Notably, for QVHighlights, this improvement is less pronounced. We hypothesize this is due to the videos in this benchmark having constant duration. Nevertheless, it still performs far worse than the Chrono-BLIP model.
Compared to prior SOTA methods on QVHighlights (QVH), Chrono-BLIP outperforms the previous SOTA InternVideo2 [Wang2024InternVideo2SF] on all metrics.
On Charades-STA, the Chrono-BLIP model also achieves a new state-of-the-art, improving over the previous SOTA model [Wang2024InternVideo2SF] for R1@0.7. While InternVideo2 achieves good performance by extracting intermediate features and leveraging CG-DETR [moon2023CQ_DETR] as a localization head, we demonstrate strong results with a fraction of the training data, smaller model size, fewer trained parameters, and while not relying on a task-specific output head. Chrono-BLIP also outperforms the previous SOTA [yan2023unloc] in moment retrieval on ActivityNet Captions by 5.62% and 5.35% on R1@0.5 and R1@0.7, respectively.
The Chrono model achieves performance that is 20.14% and 23.29% higher than SeViLa on QVHighlights in Recall@1 at thresholds of IoU=0.5 and IoU=0.7, while using the same backbone, BLIP-2 [li2023blip2]. This demonstrates the effectiveness of the novel multimodal sequence design. In contrast, SeViLa classifies each frame as relevant to the query or not without any contextual and temporal information about the video.
The results for Chrono-Qwen follow similar trends to those of Chrono-BLIP, indicating that the Chrono blueprint generalizes beyond a single backbone to decoder-only architectures of different sizes and training recipes.
Moreover, we also evaluate the Chrono-GPT model on all three MR benchmarks in a zero-shot setting. As expected, the zero-shot performance achieved by Chrono-GPT is lower compared to methods explicitly trained for Moment Retrieval. However, Chrono-GPT performs significantly better than the baseline GPT-4o. We observe that without any training or timestamps, GPT-4o completely fails to associate specific frames with a specific time, even when provided with the total duration and the information that the frames are sampled uniformly. We suspect the temporal encoding and training task distribution significantly affect this ability. Therefore, in Table VII, we show results with Qwen-2.5-VL models in the zero-shot setting, which employ a positional encoding method that explicitly represents the time associated with each frame.
Prior work [qian2024momentor, Ren2023TimeChatAT, huang2024vtimellm] performs extensive video-language instruction finetuning on multiple temporal grounding datasets and then evaluates on a held-out target dataset (commonly Charades-STA). To compare under a clear protocol, we report cross-dataset transfer results in Table II, which we define as finetuning on a source dataset and evaluating on a different target dataset without any further tuning on the target. Concretely, we use two pairs: ActivityNet Captions $\rightarrow$ Charades-STA, and QVHighlights $\rightarrow$ ActivityNet Captions, keeping the same training setup and frame counts as in the main experiments. Despite training on a single source dataset (versus broad instruction-tuning in prior work), Chrono-BLIP achieves competitive or superior transfer performance.
This demonstrates the surprising effectiveness of simple but deliberate design choices for leveraging an image-text pretrained MLLM to achieve state-of-the-art results.
TABLE II: Transfer Learning Results for moment retrieval across datasets. indicates that Momentor was trained on a subset of ActivityNet Captions.
Charades-STA ActivityNet
Method R1@.5 R1@.7 R1@.5 R1@.7
Momentor [qian2024momentor] 26.60 11.60 23.00 12.40
TimeChat [Ren2023TimeChatAT] 32.20 13.40
VTimeLLM [huang2024vtimellm] 34.30 14.70
Chrono-BLIP[Uncaptioned image] 34.87 18.04 28.64 16.44
4.4 Grounded Video QA on NExT-GQA
Next, we challenge Chrono to a new task, grounded video question answering (GQA) on the NExT-GQA [Xiao2023NExT-QGA] benchmark, i.e., localizing a moment that is relevant to answering a question and, then even answering the question.
4.4.1 Localize-then-answer using Chrono-BLIP
To evaluate Chrono-BLIP on GQA, we follow the localizer-answerer approach of SeViLa [yu2023SeViLA] using 60 frames for localizing the relevant moment, and then resampling 60 new frames out of that segment for answering the question. We apply the QVH-pretrained Chrono-BLIP model as the localizer and finetune a separate Flan-T5 XL LLM as the answerer. Notably, we do not finetune the localization model, thus applying it to a new domain, the domain of questions, where a relevant moment does not have to be explicitly described in the question to be relevant to answering it. We also compare to uniformly sampling 60 frames. Table I(b) demonstrates the performance of Chrono-BLIP on the NExT-GQA [Xiao2023NExT-QGA] benchmark and compares it to prior state-of-the-art models, SeViLa and FrozenBiLM (NG+) [Xiao2023NExT-QGA].
4.4.2 Single localizer-answerer using Chrono-GPT
Following the same protocol, we evaluate the Chrono-GPT model in a zero-shot setting. The timestamp design enables strong zero-shot moment retrieval for Chrono-GPT (see Section 4.6 for the ablations). Paired with the strong QA abilities of a foundational model, we achieve a new state-of-the-art result on NExT-GQA over prior work [Zhang2023LLoVi, Qin2024DeVi], which also leverages GPT-4 and GPT-4o, respectively.
The best Chrono-GPT setup for NExT-GQA uses a single stage, i.e., returns answer and temporal grounding in the same model response. We find that this performs better than 2 stages, where the model first performs moment retrieval, and then the model is asked to answer the question. This is the case even when the second stage is able to zoom in on the window, use more frames overall, and keeps the first stage in context. We choose to use 1.4FPS for the final results. This offers 1-5% relative improvement in all metrics compared to using 60 frames for every video. This uses an equivalent number of frames across the entire dataset (with an average length of around 40 seconds) compared to using 60 frames for every video. Using absolute interleaved timestamps significantly improves temporal grounding and, therefore, grounded accuracy. We find that grounding modestly helps overall accuracy, i.e., not thresholded by grounding, as can be seen also in Table III. This could indicate that including temporal grounding capabilities into models that process video cannot only offer greater interpretability but also improve their overall performance at other tasks like video QA. Additionally, this indicates that the increase in grounded accuracy is not strongly driven by better answering capabilities, but truly better temporal localization.
TABLE III: Does moment retrieval help Video QA? Asking the model to ground its answer provides modest improvements in accuracy, with grounding accuracy significantly increasing when using Chrono[Uncaptioned image]. For cost reasons, we only perform this ablation on 1000 samples from NExT-GQA validation split, using GPT-4o.
MR Chrono[Uncaptioned image] mIoU mIoP IoP@.5 A@GQA A@QA
- - - - 79.80
22.47 29.85 27.13 22.03 80.07
37.02 51.62 52.45 43.35 80.15
4.5 Qualitative Results
Figure 3 shows qualitative results that demonstrate Chrono-BLIP’s ability to predict multiple windows (2 and 3), generalize to unseen data distributions (5), as well as some failure modes like predicting wrong segments (3) and only partially adhering to the prompt due to low image resolution (4). In (4), the query is ”A woman in long brown hair is trying on a black hat in a shop”. Chrono-BLIP predicts a moment where the woman is clearly visible in a shop but not trying on a hat. In the ground truth segment, the woman is trying on a hat but is not clearly visible since she is far away. We provide more videos with localizations in Appendix C.
4.6 Ablation Studies
By default, the Chrono models for Charades-STA (QVHighlights / ActivityNet Captions) leverage 20 (60) visual frames, their associated timestamps, video duration, a query, and a task prompt. If not stated otherwise, we represent time as simple text tokens of whole seconds and interleave these with frame tokens. In the following, we ablate the impact of these design choices on the downstream moment retrieval performance by reporting results for Chrono-BLIP on the Charades-STA validation set and for Chrono-GPT on the QVHighlights validation set. Moreover, we provide further evidence for the generalization of the blueprint by performing ablations on the Qwen2.5-VL [Qwen2.5-VL] family of models across different model sizes in zero-shot and finetuned settings. In several of the ablation experiments, we also report mIoU for completeness, although we find the relative performance in terms of mIoU follow those of R1@0.5, which is, in any case, more informative of demanding temporal grounding criteria. Therefore, in many other cases we report only R1@0.5 and R1@0.7.
@@ 4.6 Ablation Studies
TABLE IV: Ablation studies on input sequence and timestamp design. For finetuned Chrono-BLIP, we ablate on Charades-STA, and for zero-shot Chrono-GPT, on QVHighlights. We use validation set in all cases. D: Video Duration, T: Frame Timestamps, Rep: Representation, Rel: Relative, Abs: Absolute, Prec: Precision, Dec: Decimal, Int: Integer, Inter: Interleaved. Bold indicates best result for a given base model and metric. Underlined indicates second best.
@@
4.6.1 Input sequence design
In Table IV (a), we analyze the effectiveness of each part of the multimodal input sequence of the MLLM by cumulatively adding each component, starting with only providing the frames. The query is always part of the sequence. For Chrono-BLIP, adding the video duration, in addition to the visual inputs, improves the downstream performance significantly (row 2 vs. row 1). This shows the importance of providing the model a point of reference to infer the temporal location of each frame and demonstrates the ability of the MLLM to learn this association. ChronoGPT, on the other hand, does not benefit as strongly from the additional information as the fine-tuned BLIP-2 model. Again, note that with a fixed number of uniformly sampled frames and the video duration, the model has all the information necessary to compute at which timestamp each frame was sampled. This experiment shows that GPT-4o appears not to be able to leverage this information.
Yet, providing the specific timestamps at which each frame was sampled in an interleaved manner (as illustrated in Figure 2) improves the model performance of both models by a significant margin (rows (3, 4) vs. rows (1,2)). This demonstrates the relevance of designing the input context to provide as much relevant information as possible while still not relying on additional computation, such as generating and leveraging transcripts. Providing all timestamps along with the video duration yields the best performance (row 4 vs. 3) for both models.
TABLE V: Input components for zero-shot moment retrieval using GPT-4o on Charades-STA and QVHighlights val sets.
Charades-STA QVHighlights
Input R1@.5 R1@.7 mIoU R1@.5 R1@.7 mIoU
Frames only $8.62$ $2.67$ $14.74$ $5.49$ $2.17$ $10.56$
+ Duration $11.56$ $4.59$ $18.25$ $4.32$ $1.87$ $8.96$
+ Timestamps 28.76 10.99 35.43 61.68 41.80 57.12
Complementary to Table IV (a), Table V isolates the zero-shot GPT-4o setting on Charades-STA and QVHighlights. Starting from frames only, adding the video duration yields only minor gains, whereas including interleaved timestamps in Chrono-GPT[Uncaptioned image] leads to large improvements across R1@0.5/0.7 and mIoU. This mirrors the behavior observed for the finetuned Chrono-BLIP model, reinforcing that the interleaved timestamp design is crucial even for powerful zero-shot MLLMs such as GPT-4o.